a body thrown up with finite speed is caught back after 4sec

Grade:11

2 Answers

Arpit kumar verma
7 years ago
Please give full details about question. What i can find out?Now if u ask for velocity then velocity is 2g m/sec
Aman Verma
5 years ago
At highest point  V=0
t=4s/2=2sec
Now use 
V=u+gt
0=u-10*2(g=-10,t=10s
u=20m/s